A deal has two kinds of mets; one costs 90 cents a pound, the other 60 cents a pound. He wishes to make 50 pounds of a mixture that will cost 72 cents a pound. How many pounds of each kind should he use?
Answer by checkley71(8403)

.90X+.60(50-X)=.72*50
.90X+30-.60X=36
.30X=36-30
.30X=6
X=6/.30
X=20 AMOUNT OF $.90 MEAT.
50-20=30 AMOUNR OF $.60 MEAT.
PROOF:
.90*20+.60*30=36
18+18=36
36=36